OpenStudy (anonymous):

Give the 3rd and 7th terms using the sequence formula tn = (n – 3)2. 0, 16 0, 8 0, 49 0, 14

OpenStudy (anonymous):

replace \(n\) by 3 get \((3-3)^2=0\)

OpenStudy (anonymous):

replace \(n\) by 7 get \((7-3)^2=?\)

OpenStudy (anonymous):

49-9 = 40 so shouldnt the answer be (0,40)?

OpenStudy (anonymous):

thats why i isss so confused